Grade 11Physical Chemistry

How much copper can be obtained from 100 g of copper sulphate?

Sumeet phangso
9 Years agoGrade 11

1 Answer

Niraj
9 Years ago
Cu=64 u(approx)160 g(GMM) of CuSO4 contains 64 g of Cu.1g of CuSO4=64/160g of Cu.Therefore, 100g of CuSO4=(64*100)/160. (Calculate) 100g of CuSO4=40g of Cu.